Complete and balance the following acid-base and acid-carbonate reactions:
HCl+Ba(OH)2=
Hcl+Al(CO3)3=
H3PO4+LiHCO3=
Al(OH)3+H2SO4=
1
Expert's answer
2015-12-01T11:03:20-0500
2HCl + Ba(OH)2 = BaCl2 + 2H2O
6HCl + Al2(CO3 )3 = 2AlCl3 + 3H2O + 3CO2
H3PO4 + 3LiHCO3 = Li3 PO4 + 3CO2 + 3H2O
2Al(OH)3 + 3H2SO4 = Al2(SO4)3 + 6H2O